The function y=60/x relates y, the time in hours that it takes Michael to travel to his grandparents’ house, to x, his average speed in miles per hour. What does the number 60 in the function represent?
Found 2 solutions by jorel555, Alan3354:
Answer by jorel555(1290)

60 represents the number of minutes in an hour!!!!!!!!!!!!!!

Answer by Alan3354(69443)

The function y=60/x relates y, the time in hours that it takes Michael to travel to his grandparents’ house, to x, his average speed in miles per hour. What does the number 60 in the function represent?
--------------
y = 60/x
x*y = 60
(mi/hr)*hr = ??